a baseball is hit so that it travels straight upward after being struck by the bat. a fan observes that it takes 3.00 seconds to reach it's maximum height. find (A) the ball initial velocity and (B) the height it reaches.
1
Expert's answer
2013-03-14T10:31:55-0400
a baseball is hit so that it travels straight upward after beingstruck by the bat. a fan observes that it takes 3.00 seconds to reach it's
maximum height. find

(A) the ball initialvelocity
The ball falls freelly,therefore:
v = v0 - g*t
v0 - the ball initialvelocity
g - gravitationalacceleration
t - time
When v = 0, t = 3.00sec:
0 = v0 - g*t => v0 = g*t
v0 = 9.8*3m/s= 29.4 m/s
Answer: v0 = 29.4 m/s

(B) the height itreaches.
The law conservation ofenergy:
m*g*h + 0= m*v0^2/2 + 0
h - maximum height
h = v0^2/(2g) = | v0 = g*t | = gt^2/2 = 44.1 m
Answer: h = 44.1 m